# FLIPHAT: Joint Differential Privacy for High Dimensional Sparse Linear Bandits

### v2

@@ Appendix E Proofs in Section 5
For control on $\left\|\tilde{\mathbf{w}}_{S}\right\|_{2}^{2}$, we note that in our case $\lambda=\eta BT\sqrt{3s\log(M_{\ell}/\delta)}/(N\epsilon)$, with $B\asymp R_{\ell}$, $s\asymp s^{*}$ and $M_{\ell}=28\kappa\log(2b_{\sf max}^{2}N_{\ell})\asymp\log N_{\ell}$, which gives
$\lambda\asymp(\log N_{\ell})^{2}(s^{*})^{1/2}\sqrt{\log((\log N_{\ell})/\delta)}/N_{\ell}\epsilon.$
Using (17) from the above lemma with $k=s\asymp s^{*}$, $\lambda$ as above and $C=c_{1}^{\prime}\log d$ so that
$kC^{2}\lambda^{2}=s^{*}\times 16\log^{2}d\times\frac{s^{*}\log((\log N_{\ell})/\delta)\log^{3}N_{\ell}}{N_{\ell}^{2}\epsilon^{2}}\asymp\frac{(s^{*})^{2}\log^{2}d\log((\log N_{\ell})/\delta)\log^{2}N_{\ell}}{N_{\ell}^{2}\epsilon^{2}}$
@@
${\mathbb{P}}\left(\left\|\tilde{\mathbf{w}}_{S}\right\|_{2}^{2}>K.\frac{(s^{*})^{2}\log^{2}d\log((\log N_{\ell})/\delta)\log^{3}N_{\ell}}{N_{\ell}^{2}\epsilon^{2}}\right)\leq\frac{s^{*}}{d^{8}}.$
@@
$C^{2}\lambda^{2}\log^{2}k=C^{2}\times\frac{s^{*}\log((\log N_{\ell})/\delta)\log^{3}N_{\ell}}{N_{\ell}^{2}\epsilon^{2}}\times\log^{2}d\asymp\frac{(s^{*})\log^{2}d\log((\log N_{\ell})/\delta)\log^{3}N_{\ell}}{N_{\ell}^{2}\epsilon^{2}}$
@@
$P\left(\left\|\mathbf{w}_{i}\right\|_{\infty}^{2}>K^{\prime}.\frac{(s^{*})\log^{2}d\log((\log N_{\ell})/\delta)\log^{3}N_{\ell}}{N_{\ell}^{2}\epsilon^{2}}\right)\leq\frac{1}{d^{8}}.$
@@
$P\left(\mathbf{W}_{m}\leq K^{\prime}.\frac{(s^{*})^{2}\log^{2}d\log((\log N_{\ell})/\delta)\log^{3}N_{\ell}}{N_{\ell}^{2}\epsilon^{2}}\right)\geq\left(1-\frac{s^{*}}{d^{8}}\right)\left(1-\frac{1}{d^{8}}\right)^{s}\geq 1-\frac{s+s^{*}}{d^{8}}.$
@@
$P\left(\max_{m\in[M_{\ell}]}\mathbf{W}_{m}>K^{\prime}.\frac{(s^{*})^{2}\log^{2}d\log((\log N_{\ell})/\delta)\log^{3}N_{\ell}}{N_{\ell}^{2}\epsilon^{2}}\right)$
$\leq\sum_{m}P\left(\mathbf{W}_{m}>K^{\prime}.\frac{(s^{*})^{2}\log^{2}d\log((\log N_{\ell})/\delta)\log^{3}N_{\ell}}{N_{\ell}^{2}\epsilon^{2}}\right)$
$\leq\frac{M_{\ell}(s+s^{*})}{d^{8}}=O(1/d^{6})$
where the last line follows from $M_{\ell}\asymp N_{\ell}\lesssim d$ and $s\asymp s^{*}\lesssim d$.
@@ E.3 Proof of Regret Bound
$\lesssim 2x_{\max}b_{\max}(2^{L}-1)+2x_{\max}\sum_{\ell\geq L}(t_{\ell+1}-t_{\ell})\sigma^{2}\kappa_{0}\left(\sqrt{\frac{{s^{*}}^{2}\log d}{N_{\ell-1}}}+\sqrt{\frac{{s^{*}}^{3}(\log d)^{2}\log(\log N_{\ell-1}/\delta)\log^{3}N_{\ell-1}}{N_{\ell-1}^{2}\epsilon^{2}}}\right)+O(1)$
$\leq x_{\max}b_{\max}(2^{L}-1)+4\sigma\kappa_{0}x_{\max}\sum_{\ell\geq L}\left[{s^{*}}\sqrt{N_{\ell-1}\log d}+\sqrt{\frac{{s^{*}}^{3}(\log d)^{2}\log(1/\delta)\log^{4}N_{\ell-1}}{\epsilon^{2}}}\right]$
$\leq x_{\max}b_{\max}(2^{L}-1)+4\sigma^{2}\kappa_{0}x_{\max}{s^{*}}\sqrt{\log d}\sum_{\ell\geq L}2^{(\ell-1)/2}+4\sigma\kappa_{0}x_{\max}\sqrt{\frac{{s^{*}}^{3}(\log d)^{2}\log(1/\delta)}{\epsilon^{2}}}\sum_{\ell\geq L}(\ell-1)^{2}$
$\lesssim x_{\max}b_{\max}(2^{L}-1)+4\sigma\kappa_{0}x_{\max}{s^{*}}\sqrt{T\log d}+4\sigma\kappa_{0}x_{\sf max}\sqrt{\frac{{s^{*}}^{3}(\log d)^{2}\log(1/\delta)\log^{6}T}{\epsilon^{2}}}$
$\lesssim 4\sigma\kappa_{0}x_{\max}{s^{*}}\sqrt{T\log d}+4\sigma\kappa_{0}x_{\sf max}\sqrt{\frac{{s^{*}}^{3}(\log d)^{2}\log(1/\delta)\log^{6}T}{\epsilon^{2}}}.$
Margin bound
@@
$\phi_{\ell(t)}\asymp x_{\sf max}\sigma^{2}\kappa_{0}\left(\sqrt{\frac{{s^{*}}^{2}\log d}{N_{\ell-1}}}+\sqrt{\frac{{s^{*}}^{3}(\log d)^{2}\log(1/\delta)\log^{4}N_{\ell-1}}{N_{\ell-1}^{2}\epsilon^{2}}}\right).$
@@
$24x_{\sf max}\sigma^{2}\kappa_{0}\underbrace{\sum_{\ell\geq L}\Delta_{*}^{-\alpha}N_{\ell-1}\left[\left(\frac{{s^{*}}^{2}\log d}{N_{\ell-1}}\right)^{(1+\alpha)/2}+\left(\frac{{s^{*}}^{3}(\log d)^{2}\log(1/\delta)\log^{4}N_{\ell-1}}{N_{\ell-1}^{2}\epsilon^{2}}\right)^{(1+\alpha)/2}\right]}_{I_{\alpha}}+O(1).$
@@
$I_{\alpha}\lesssim\frac{{s^{*}}^{1+\alpha}(\log d)^{(1+\alpha)/2}}{\Delta_{*}^{\alpha}}\left(\frac{T^{\frac{1-\alpha}{2}}-1}{1-\alpha}\right)+\Psi_{\alpha}\Delta_{*}^{-\alpha}\epsilon^{-(1+\alpha)}\{\log(1/\delta)\}^{(1+\alpha)/2}\left({s^{*}}^{3}\log^{2}d\right)^{\frac{1+\alpha}{2}},$
@@
$I_{1}\lesssim\frac{{s^{*}}^{2}\log d}{\Delta_{*}}\log T+\frac{\Psi_{1}}{\Delta_{*}\epsilon^{2}}\log(1/\delta){s^{*}}^{3}\log^{2}d.$
@@
$I_{\alpha}\lesssim\frac{{s^{*}}^{1+\alpha}(\log d)^{(1+\alpha)/2}}{\Delta_{*}^{\alpha}}\left(\frac{1-T^{-\frac{\alpha-1}{2}}}{\alpha-1}\right)+\Psi_{\alpha}\Delta_{*}^{-\alpha}\epsilon^{-(1+\alpha)}\{\log(1/\delta)\}^{(1+\alpha)/2}\left({s^{*}}^{3}\log^{2}d\right)^{\frac{1+\alpha}{2}},$
@@
$\ell(t)\gtrsim{s^{*}}^{2}\log d+\frac{{s^{*}}^{3}(\log d)^{2}\log(1/\delta)}{\epsilon^{2}}.$

### v3

@@ Appendix E PROOFS IN SECTION 5
For control on $\left\|\tilde{\mathbf{w}}_{S}\right\|_{2}^{2}$, we note that in our case $\lambda=\eta BT\sqrt{3s\log(M_{\ell}/\delta)}/(N\epsilon)$, with $B\asymp R_{\ell}$ and $M_{\ell}=28\kappa\log(2b_{\sf max}^{2}N_{\ell})\asymp\log N_{\ell}$, which gives
$\lambda\asymp(\log N_{\ell})^{2}s^{1/2}\sqrt{\log((\log N_{\ell})/\delta)}/N_{\ell}\epsilon.$
Using (18) from the above lemma with $k=s$, $\lambda$ as above and $C=c_{1}^{\prime}\log d$ so that
$kC^{2}\lambda^{2}=s\times 16\log^{2}d\times\frac{s\log((\log N_{\ell})/\delta)\log^{3}N_{\ell}}{N_{\ell}^{2}\epsilon^{2}}\asymp\frac{s^{2}\log^{2}d\log((\log N_{\ell})/\delta)\log^{2}N_{\ell}}{N_{\ell}^{2}\epsilon^{2}}$
@@
${\mathbb{P}}\left(\left\|\tilde{\mathbf{w}}_{S}\right\|_{2}^{2}>K^{\prime}.\frac{s^{2}\log^{2}d\log((\log N_{\ell})/\delta)\log^{3}N_{\ell}}{N_{\ell}^{2}\epsilon^{2}}\right)\leq\frac{s}{d^{8}}.$
@@
$C^{2}\lambda^{2}\log^{2}k=C^{2}\times\frac{s\log((\log N_{\ell})/\delta)\log^{3}N_{\ell}}{N_{\ell}^{2}\epsilon^{2}}\times\log^{2}d\asymp\frac{s\log^{2}d\log((\log N_{\ell})/\delta)\log^{3}N_{\ell}}{N_{\ell}^{2}\epsilon^{2}}$
@@
$P\left(\left\|\mathbf{w}_{i}\right\|_{\infty}^{2}>K^{\prime}.\frac{s\log^{2}d\log((\log N_{\ell})/\delta)\log^{3}N_{\ell}}{N_{\ell}^{2}\epsilon^{2}}\right)\leq\frac{1}{d^{8}}.$
@@
$P\left(\mathbf{W}_{m}\leq K^{\prime}.\frac{s^{2}\log^{2}d\log((\log N_{\ell})/\delta)\log^{3}N_{\ell}}{N_{\ell}^{2}\epsilon^{2}}\right)\geq\left(1-\frac{s}{d^{8}}\right)\left(1-\frac{1}{d^{8}}\right)^{s}\geq 1-\frac{2s}{d^{8}}.$
@@
$P\left(\max_{m\in[M_{\ell}]}\mathbf{W}_{m}>K^{\prime}.\frac{s^{2}\log^{2}d\log((\log N_{\ell})/\delta)\log^{3}N_{\ell}}{N_{\ell}^{2}\epsilon^{2}}\right)$
$\leq\sum_{m}P\left(\mathbf{W}_{m}>K^{\prime}.\frac{s^{2}\log^{2}d\log((\log N_{\ell})/\delta)\log^{3}N_{\ell}}{N_{\ell}^{2}\epsilon^{2}}\right)$
$\leq\frac{M_{\ell}\times 2s}{d^{8}}=O(1/d^{6})$
where the last line follows from $M_{\ell}\asymp N_{\ell}\lesssim d$ and $s\lesssim d$.
@@ E.3 Proof of Regret Bound
$\lesssim 2x_{\max}b_{\max}(2^{L}-1)+2x_{\max}\sum_{\ell\geq L}(t_{\ell+1}-t_{\ell})\sigma^{2}\kappa_{0}\left(\sqrt{\frac{(s+s^{*})^{2}\log d}{N_{\ell-1}}}+\sqrt{\frac{(s+s^{*})^{3}(\log d)^{2}\log(\log N_{\ell-1}/\delta)\log^{3}N_{\ell-1}}{N_{\ell-1}^{2}\epsilon^{2}}}\right)$
$\quad+O(1)$
$\leq x_{\max}b_{\max}(2^{L}-1)+4\sigma\kappa_{0}x_{\max}\sum_{\ell\geq L}\left[(s+s^{*})\sqrt{N_{\ell-1}\log d}+\sqrt{\frac{(s+s^{*})^{3}(\log d)^{2}\log(1/\delta)\log^{4}N_{\ell-1}}{\epsilon^{2}}}\right]$
$\leq x_{\max}b_{\max}(2^{L}-1)+4\sigma^{2}\kappa_{0}x_{\max}(s+s^{*})\sqrt{\log d}\sum_{\ell\geq L}2^{(\ell-1)/2}+4\sigma\kappa_{0}x_{\max}\sqrt{\frac{(s+s^{*})^{3}(\log d)^{2}\log(1/\delta)}{\epsilon^{2}}}\sum_{\ell\geq L}(\ell-1)^{2}$
$\lesssim x_{\max}b_{\max}(2^{L}-1)+4\sigma\kappa_{0}x_{\max}(s+s^{*})\sqrt{T\log d}+4\sigma\kappa_{0}x_{\sf max}\sqrt{\frac{(s+s^{*})^{3}(\log d)^{2}\log(1/\delta)\log^{6}T}{\epsilon^{2}}}$
$\lesssim 4\sigma\kappa_{0}x_{\max}(s+s^{*})\sqrt{T\log d}+4\sigma\kappa_{0}x_{\sf max}\sqrt{\frac{(s+s^{*})^{3}(\log d)^{2}\log(1/\delta)\log^{6}T}{\epsilon^{2}}}.$
E.3.1 Margin Bound
@@
$\phi_{\ell(t)}\asymp x_{\sf max}\sigma^{2}\kappa_{0}\left(\sqrt{\frac{(s+s^{*})^{2}\log d}{N_{\ell-1}}}+\sqrt{\frac{(s+s^{*})^{3}(\log d)^{2}\log(1/\delta)\log^{4}N_{\ell-1}}{N_{\ell-1}^{2}\epsilon^{2}}}\right).$
@@
$24x_{\sf max}\sigma^{2}\kappa_{0}\underbrace{\sum_{\ell\geq L}\Delta_{*}^{-\alpha}N_{\ell-1}\left[\left(\frac{(s+s^{*})^{2}\log d}{N_{\ell-1}}\right)^{(1+\alpha)/2}+\left(\frac{(s+s^{*})^{3}(\log d)^{2}\log(1/\delta)\log^{4}N_{\ell-1}}{N_{\ell-1}^{2}\epsilon^{2}}\right)^{(1+\alpha)/2}\right]}_{I_{\alpha}}+O(1).$
@@
$I_{\alpha}\lesssim\frac{(s+s^{*})^{1+\alpha}(\log d)^{(1+\alpha)/2}}{\Delta_{*}^{\alpha}}\left(\frac{T^{\frac{1-\alpha}{2}}-1}{1-\alpha}\right)+\Psi_{\alpha}\Delta_{*}^{-\alpha}\epsilon^{-(1+\alpha)}\{\log(1/\delta)\}^{(1+\alpha)/2}\left((s+s^{*})^{3}\log^{2}d\right)^{\frac{1+\alpha}{2}},$
@@
$I_{1}\lesssim\frac{(s+s^{*})^{2}\log d}{\Delta_{*}}\log T+\frac{\Psi_{1}}{\Delta_{*}\epsilon^{2}}\log(1/\delta)(s+s^{*})^{3}\log^{2}d.$
@@
$I_{\alpha}\lesssim\frac{(s+s^{*})^{1+\alpha}(\log d)^{(1+\alpha)/2}}{\Delta_{*}^{\alpha}}\left(\frac{1-T^{-\frac{\alpha-1}{2}}}{\alpha-1}\right)+\Psi_{\alpha}\Delta_{*}^{-\alpha}\epsilon^{-(1+\alpha)}\{\log(1/\delta)\}^{(1+\alpha)/2}\left((s+s^{*})^{3}\log^{2}d\right)^{\frac{1+\alpha}{2}},$
@@
$\ell(t)\gtrsim(s+s^{*})^{2}\log d+\frac{(s+s^{*})^{3}(\log d)^{2}\log(1/\delta)}{\epsilon^{2}}.$
@@
Appendix F NUMERICAL EXPERIMENTS: MORE DETAILS
We discuss the numerical experiments in further details and also provide additional experiments for (i) different noise distribution, (ii) effect of the tuning parameter $s$ and (iii) effect of the number of arms $K$. For implementation of FLIPHAT, we used three hyperparameters $s$ (sparsity guess), $\eta$ (step size for N-IHT) and $M>0$ (to control the number of steps in N-IHT with $M_{\ell}=M\log N_{\ell}$ at the $\ell$-th episode). For these synthetic experiments, we used $x_{\sf max}$ and $b_{\sf max}$ based on the true underlying setting and chose $C,R_{\ell}$ accordingly (see Algorithm 2 and Proposition 5.5. In practice, the contexts can be rescaled (to eliminate $x_{\sf max}$) and $b_{\sf max}$ could either be chosen based on existing domain knowledge or held-out data or included as another hyperparameter. The task of tuning such hyperparameters can be tricky because of privacy constraints, but we do not focus on that aspect in this paper and keep it for future work.
Figure 2: Regret $R(T)$ versus time horizon $T$ in the setting $d=400,s^{*}=5,K=3$, with different privacy parameters $\epsilon$ and $\delta=0.01$. Contexts iid from normal with autoregressive covariance. (Left) Gaussian observation noise, mean 0 and $\sigma=0.1$, (Right) Bounded observation noise, drawn from $\text{Uniform}(-0.1,0.1)$
F.1 Regret vs $T$
We perform two experiments to study how the regret $R(T)$ depends on time horizon $T$ for different levels of the privacy parameter $\epsilon$. For both experiments, we consider $d=400,s^{*}=5,K=3$ and choose the privacy parameters $\epsilon\in\{0.5,1,2,5,10\}$ and $\delta=0.01.$ The contexts were drawn i.i.d. from $\mathcal{N}(0,\Sigma)$ with an auto-regressive covariance matrix $\Sigma_{ij}=0.1^{|i-j|}$. The rewards were sampled from the linear model (1) – in the first experiment $\epsilon(t)\sim\mathcal{N}(0,\sigma)$ with $\sigma=0.1$ and the second experiment, $\epsilon(t)\sim\text{Uniform}(-0.1,0.1)$. The hyperparameters were chosen as $s=10,\eta=10^{-4},M=1.6$. Each experiment was repeated 60 times and the mean and 95% confidence intervals are shown in Figure 2. We find the behavior very similar in both cases. For additional perspective, we added the random bandit (full privacy $\epsilon=0$) and sparsity agnostic bandit (no privacy $\epsilon\to\infty$). We note that even for $\epsilon=0.5$, we observe clear sub-linear behavior of the regret, particularly after $T=7500$. For larger $\epsilon$, we observe that the performance is similar to the sparsity agnostic bandit. To our knowledge, this current paper is the first work on joint differentially private algorithm for sparse contextual bandits, as a result, we could not find any other suitable baseline to compare the performance with.
F.2 Regret vs $d$
The setup is exactly similar to the first two experiments, discussed above. In this case, we only collect the regret suffered at $T=10000$, but use different context dimensions $d$, to examine the effect of $d$ on the regret for different choices of $\epsilon$. Note that $s^{*}$ is fixed in these experiments. The result is shown in Figure 1 (right) in the main text. We wish to remark on the distinct behavior of FLIPHAT, compared to the JDP contextual bandit in (Shariff and Sheffet,, 2018). In this work, the authors explore joint differential privacy for contextual bandits, with no additional condition on the sparsity – as a result, the regret scales as $d^{2}$, which is illustrated in Figure 3 (in the Supplementary Material of (Shariff and Sheffet,, 2018)). In our case, we demonstrate that with the additional assumption of sparsity, we can improve this dependence on $d$ under joint differential privacy – this is illustrated in Figure 1 (right), where we observe that $R_{d}(T)\approx\alpha+\beta\log d$, where $\beta$ (for different $\epsilon$) is estimated using a linear model and shown in the legend.
F.3 Effect of $s$
In this experiment, the goal was to examine the effect of one of the more important hyperparameters $s$. According to Theorem 5.6, as long as $s\gtrsim\kappa^{2}s^{*}$, the upper bound (in the problem independent setting of $\alpha=0$), scales as $(s+s^{*})^{3/2}$. As discussed in the remarks following the theorem, setting $s$ in the order of $o(d)$ guarantees improved dependence on $d$, with $s\asymp s^{*}$ (same order) yielding near-optimal dependence on $s^{*}$. In practice, $s^{*}$ is not known, although some domain knowledge might give reasonable lower (and upper) bounds for it. Figure 3 (left) shows $\log R(T)$ against $\log s$ for $T=10000$ in the same setting as the first experiment (with the only change being $s^{*}=10$ here) for different choices of $s$ ranging from 5 to 60 (and different $\epsilon$). We observe that the regret is poor for lower values and is best around $s=s^{*}$ and then worsens a bit. The performance is reasonably robust for higher values of $s$ under high privacy (small $\epsilon$) regime. Under low privacy regime, the log-regret scales only linearly in $\log s$. The slope of the $\log R(T)$ vs $\log s$ lines for each $\epsilon$ in the part $s\geq s^{*}$ is estimated by fitting a linear model and displayed in the corresponding legend.
F.4 Effect of $K$
In our analysis, we held the number of arms $K$ fixed (small compared to $d,T$). We conduct numerical experiment to see the effect of $K$ on $R(T)$ for FLIPHAT with $T=10000$, in the setting of the first experiment $d=400,s^{*}=5$, at different levels of privacy. The results are shown in Figure 3 (right), where we notice that the dependence is indeed sub-linear.
Figure 3: (Left) Effect of the choice of the tuning parameter $s$ on regret $R(T)$, with $T=10000$ true $s^{*}=10$ for different privacy levels. (Right) Regret at $T=10000$ for different number of arms $K$, keeping $d,s^{*}$ and other parameters fixed.